# Clutter-Aware Target Detection for ISAC in a Millimeter-Wave Cell-Free Massive MIMO System

### v2

@@ III-A Power allocation optimization
In this subsection, we introduce the first subproblem of the AO algorithm, where we optimize the power allocation for a fixed RIS configuration. We first reformulate (12) as
$\overline{\textrm{SCNR}}=\boldsymbol{\rho}^{\top}\Re(\mathbf{A})\boldsymbol{\rho},$ (14)
where
@@
The set of constraints (13b) can be reformulated as second order cone (SOC) constraints in terms of $\boldsymbol{\rho}$ as
$\left\lVert\begin{bmatrix}|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{1}|\sqrt{\rho_{1}}&\dots&\underbrace{0}_{k}&\dots&|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{L}|\sqrt{\rho_{L}}&\sigma_{k}\end{bmatrix}\right\rVert$
$\leq\frac{|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{k}|\sqrt{\rho_{k}}}{\sqrt{\gamma_{k}}}.$ (16)
The power constraint in (13c) can also be rewritten in SOC form as
$\left\lVert\text{diag}\left(\left\lVert\mathbf{f}_{1,t}\right\rVert,\dots,\left\lVert\mathbf{f}_{L,t}\right\rVert\right)\boldsymbol{\rho}\right\rVert\leq\sqrt{P_{t}}.$ (17)
The first subproblem of the AO SCNR-maximizing algorithm is defined as
@@
[…]
This problem can be solved using the convex-concave procedure outlined in [3].
@@ III-B RIS phase-shift optimization
We now move on to the second part of the AO, where we optimize $\boldsymbol{\theta}$ when the power variables are fixed. The SCNR can be reformulated as
$\overline{\textrm{SCNR}}=\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}\boldsymbol{\theta},$ (19)
where
@@
We have used $\text{Tr}(\mathbf{A}\mathbf{X}\mathbf{B}\mathbf{X}^{\mathsf{H}})=\text{vec}^{\mathsf{H}}(\mathbf{X})\left(\mathbf{B}^{\top}\otimes\mathbf{A}\right)\text{vec}(\mathbf{X})$ and the following lemma.
Lemma 1.
Let $\mathbf{x}\in\mathbb{C}^{N}$ and $\mathbf{X}=\text{diag}(\mathbf{x})$. For any $\mathbf{A}\in\mathbb{C}^{N\times N}$ and $\mathbf{B}\in\mathbb{C}^{N\times N}$, it holds that
$\text{vec}^{\mathsf{H}}(\mathbf{X})(\mathbf{A}\otimes\mathbf{B})\text{vec}(\mathbf{X})=\mathbf{x}^{\mathsf{H}}(\mathbf{A}\odot\mathbf{B})\mathbf{x}.$ (21)
Proof: Let $\text{vec}(\mathbf{X})=\sum_{n=1}^{N}x_{n}\mathbf{e}_{n}\otimes\mathbf{e}_{n}$, where $\mathbf{e}_{n}$ is the $n$-th column of $\mathbf{I}_{N}$. Then, the right-hand side of (21) can be written as $\left(\sum_{n=1}^{N}x_{n}\mathbf{e}_{n}\otimes\mathbf{e}_{n}\right)^{\mathsf{H}}(\mathbf{A}\otimes\mathbf{B})\left(\sum_{m=1}^{N}x_{m}\mathbf{e}_{m}\otimes\mathbf{e}_{m}\right)\overset{(a)}{=}\sum_{n=1}^{N}\sum_{m=1}^{N}x_{n}x_{m}^{\mathsf{H}}(\mathbf{e}_{n}^{\mathsf{H}}\mathbf{A}\mathbf{e}_{m})(\mathbf{e}_{n}^{\mathsf{H}}\mathbf{B}\mathbf{e}_{m})=\sum_{n=1}^{N}\sum_{m=1}^{N}x_{n}x_{m}^{\mathsf{H}}[\mathbf{A}]_{n,m}[\mathbf{B}]_{n,m}=\mathbf{x}^{\mathsf{H}}(\mathbf{A}\odot\mathbf{B})\mathbf{x}$, where in $(a)$ we used mixed-product property of the Kronecker product.
The RIS phase-shift optimization problem under consideration is still non-convex. We are trying to maximize a quadratic form of $\boldsymbol{\theta}$. We then employ the minorization-maximization (MM) algorithm to maximize a convex lower bound on (19) [12]. More specifically, a convex local lower bound can be found by applying the successive convex approximation [13] around the local point $\boldsymbol{\theta}^{(s)}$ at iteration $s$. The objective function can be lower bounded by
$\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}\boldsymbol{\theta}\geq 2\Re\left(\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}\right)+\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}^{(-)}\boldsymbol{\theta}-\boldsymbol{\theta}^{(s)\mathsf{H}}\mathbf{Q}^{(+)}\boldsymbol{\theta}^{(s)},$ (22)
where $\mathbf{Q}^{(+)}$ and $\mathbf{Q}^{(-)}$ represent the positive and negative semidefinite parts of $\mathbf{Q}$. We can now define the second subproblem of the AO optimization as
@@
where $\alpha$ is an auxiliary variable. Note that we have relaxed the unit modulus constraint, as it is observed that the solution satisfies this constraint with equality anyway. This is a convex problem and can be solved at each iteration until convergence is achieved. The steps of the AO algorithm are outlined in Algorithm 1.
@@ IV GLRT Detector
One of the main metrics used to assess the sensing performance is the probability of correct detection, implemented through a GLRT by extending the detector in [9] to a cell-free massive MIMO system based on the optimization procedure developed in the previous section. We operate under the assumption that the transmit signals are known by the receiving APs, thanks to the cell-free architecture. We define the vector containing the combined sensing observation of all receiving APs during timeslot $\iota$ as
$\mathbf{y}[\iota]=[\mathbf{y}_{1}^{\top}[\iota],\dots,\mathbf{y}_{R}^{\top}[\iota]]^{\top}=\mathbf{V}[\iota]\boldsymbol{\xi}+\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],$ (24)
$\mathbf{V}[\iota]=\textrm{blkdiag}(\mathbf{V}_{1}[\iota],\dots,\mathbf{V}_{R}[\iota])\in\mathbb{C}^{MR\times TR},$ (25)
$\mathbf{V}_{r}[\iota]=\left[\mathbf{E}_{1,r}\mathbf{s}_{1}[\iota],\dots,\mathbf{E}_{T,r}\mathbf{s}_{T}[\iota]\right]\in\mathbb{C}^{M\times T},$ (26)
$\mathbf{n}[\iota]=[\mathbf{n}_{1}^{\top}[\iota],\dots,\mathbf{n}_{R}^{\top}[\iota]]^{\top},\leavevmode\nobreak\ \boldsymbol{\xi}=[\xi_{1,1},\dots,\xi_{T,R}]^{\top},$ (27)
$\mathbf{U}=\textrm{blkdiag}\left(\mathbf{U}_{1},\dots,\mathbf{U}_{R}\right),\mathbf{w}[\iota]=[\mathbf{w}_{1}^{\top}[\iota],\dots,\mathbf{w}_{R}^{\top}[\iota]]^{\top}.$ (28)
@@
where $\overline{\mathbf{V}}$ is computed as $\mathbf{V}$ by using $\overline{\mathbf{E}}_{t,r}$. The adoption of a GLRT method is motivated by the unknown distribution of $\mathbf{w}[\iota]$, however we know that $\mathbf{y}_{\mathcal{H}_{0}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{MR})$ and $\mathbf{y}_{\mathcal{H}_{1}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})$, where $\mathbf{R}[\iota]=\mathbb{E}\left\{\overline{\mathbf{V}}[\iota]\boldsymbol{\xi}\boldsymbol{\xi}^{\mathsf{H}}\overline{\mathbf{V}}^{\mathsf{H}}[\iota]\right\}=\overline{\mathbf{V}}[\iota]\text{diag}(\delta^{2}_{1,1},\dots,\delta^{2}_{T,R})\overline{\mathbf{V}}^{\mathsf{H}}[\iota]$. The GLRT test is then defined as
$\frac{\prod_{\iota=1}^{\tau}\max_{\mathbf{w}[\iota]}f_{\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{1}}\left(\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{1}\right)}{\prod_{\iota=1}^{\tau}\max_{\mathbf{w}[\iota]}f_{\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{0}}\left(\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}_{0}\right)}\lessgtr^{\mathcal{H}_{0}}_{\mathcal{H}_{1}}\lambda_{d},$ (30)
where $f_{\mathbf{y}[\iota]|\mathbf{w}[\iota],\mathcal{H}}$ denote the probability density function. Following the derivation in [9], the GLRT test can be rewritten as
$\sum_{\iota=1}^{\tau}\mathbf{y}^{\mathsf{H}}[\iota]\mathbf{T}[\iota]\mathbf{y}[\iota]\lessgtr^{\mathcal{H}_{0}}_{\mathcal{H}_{1}}\underbrace{\ln(\lambda_{d})-\ln\left(\varsigma\right)}_{\widetilde{\lambda}_{d}}$ (31)
$\mathbf{T}[\iota]=\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{(}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{)}^{-1}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]$
$\quad\quad+\sigma^{-2}(\mathbf{I}_{MR}-\mathbf{U}\mathbf{U}^{\mathsf{H}})-\boldsymbol{\Xi}[\iota]$ (32)
$\boldsymbol{\Xi}[\iota]=(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})^{-1}$ (33)
where $\varsigma=\det\left(\sigma^{2}\mathbf{I}_{MR}\right)^{\tau}/{\prod_{\iota=1}^{\tau}\det\left(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR}\right)}$. $\widetilde{\lambda}_{d}$ is a revised threshold whose value should be chosen according to the desired false alarm probability, and it can be computed through Monte Carlo trials.
@@ V Numerical Results
In this section, the numerical simulation results are provided. A two-dimensional view of the simulated scenario is provided in Fig. 2. The locations of the transmit APs in $\mathcal{T}$, the receiving ones in $\mathcal{R}$, and the UEs are randomly generated within the respective bounding boxes, and their $z$ coordinates are equal to $10$ m, $10$ m, and $1$ m, respectively. The RIS and target position are assumed to be fixed and their $z$ coordinates are equal to $15$ m and $3$ m. The power budget for each transmit AP is $2$ dBW, with a bandwidth of $B=1$ MHz, the noise power is equal to $-204+10\log_{10}(B)$ dBW. The channel’s path loss follows the 3GPP urban microcell model defined in [15]. The clutter covariance matrix $\mathbf{R}_{r}$ is modeled according to the local scattering model in [16] with six clusters centered around the $\boldsymbol{\omega}_{0,r}^{\textrm{RX}}\pm[20^{\circ}$, $10^{\circ}]$. We have $T=5$ transmit APs and $R=2$ receiving APs, each equipped with $M=36$ antennas and serving $K=5$ UEs. We want to guarantee an SINR of $\gamma_{k}=3$ dB to all UEs. Unless otherwise specified, the RIS is equipped with $N=64$ elements, $C_{1}=C_{2}=C_{3}=C_{4}=2$, and the angles are uniformly distributed within an angle spread of $\sigma_{\psi_{n}}=10^{\circ}$ in azimuth and $\sigma_{\phi_{n}}=10^{\circ}$ in elevation around the LoS angles between transmitter and receiver, denoted by $\psi_{0}\leavevmode\nobreak\ (\phi_{0})$. The results shown hereafter are averaged over $10$ random realizations of the AP positions and UE positions. For each of these realizations, the probability of detection is computed over $\tau=5$ timeslots and $1000\tau$ noise and clutter realizations. The false alarm probability is set to $10^{-4}$.
@@
Figure 3: Probability of detection with respect to the RCS power with a CNR of $20$ dB.
@@
Figure 4: Probability of detection with respect to the RCS power with a CNR of $40$ dB.
@@ V-A Clutter awareness and cluster number’s impact
We first analyze the impact of the clutter awareness of the receiving APs on the detection probability and how much the LoS assumed structure of the channels affects the system’s detection performances when there are many clusters concentrated within a small angular region. Fig. 3 shows this probability of detection vs. the RCS power, where we assumed that $\delta_{1,1}^{2}=\dots=\delta_{T,R}^{2}=\delta^{2}$ for a CNR of $20$ dB. We can see that a clutter-aware system outperforms a clutter-unaware one: The clutter unaware dectector is obtained from (29) by removing the term $\mathbf{U}\mathbf{w}[\iota]$, thus obtaining $\mathbf{T}[\iota]=\sigma^{-2}\mathbf{I}_{MR}-\boldsymbol{\Xi}[\iota]$. On the other hand, when $C_{2}=C_{3}=C_{4}=20$, here denoted by $C_{234}$, the LoS assumption does not hold as much, thus generating a performance decrease. Our optimized model greatly outperforms a benchmark approach, consisting of a random RIS phase-shift matrix and an equal power allocation between the beams. Fig. 4 shows what happens when the CNR rises to $40$ dB. We see that clutter awareness rises in importance, widening the gap between the clutter-aware and unaware detectors.
@@
Figure 5: Impact of a dedicated sensing stream onto the probability of detection with a CNR of $20$ dB.
@@ V-B Do we need dedicated sensing symbols?
Given the sparse nature of the mmWave channels, one might question whether or not a dedicated sensing stream is necessary and to what extent the system benefits from it. To gain more insight into this, we have analyzed a system without sensing-dedicated precoding. Fig. 5 compares the detection performance of systems with and without dedicated sensing precoding, both in the presence and in the absence of clutter awareness. We can immediately see that removing the sensing streams implies a performance degradation. Interestingly, the system without sensing dedicated precoding seems to suffer less from the lack of clutter awareness at a low RCS value.
@@ VI Conclusions
We have investigated a cell-free massive MIMO ISAC network jointly serving multiple communication UEs while detecting a potential target through an RIS in a multi-static fashion. The system wants to maximize its sensing performance while preserving a minimum SINR for the communication UEs. To this end, we optimize the power allocation of a precoding codebook jointly with the RIS phase-shift matrix and we derive the sufficient statistics for the target detector. Numerical results show that clutter awareness at the receiving APs plays a crucial role in the system’s detection performance and that including a dedicated sensing stream benefits the system. We see that the mmWave channel structure allows the system to assume LoS channels and still achieve good performances when the channel is composed of a high number of clusters concentrated in a small angular range.

### v3

@@ III-A Power allocation optimization
In this subsection, we introduce the first subproblem of the AO algorithm, where we optimize the power allocation for a fixed RIS configuration. We first reformulate (10) as $\overline{\textrm{SCNR}}=\boldsymbol{\rho}^{\top}\Re(\mathbf{A})\boldsymbol{\rho}$ where
@@
The constraints (11b) and (11c) can be reformulated as second order cone (SOC) constraints in terms of $\boldsymbol{\rho}$. The first subproblem of the AO SCNR-maximizing algorithm is defined as
@@
subject to $\Big{[}|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{1}|\sqrt{\rho_{1}}\quad\dots\quad\underbrace{0}_{k}\quad\dots\quad|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{L}|\sqrt{\rho_{L}}$
$\sigma_{k}\Big{]}\leq(|\mathbf{g}_{k}^{\mathsf{H}}\mathbf{f}_{k}|\sqrt{\rho_{k}})/{\sqrt{\gamma_{k}}},\leavevmode\nobreak\ \forall k\in\mathcal{K}$ (13b)
$\left\lVert\widetilde{\mathbf{F}}_{t}\boldsymbol{\rho}\right\rVert\leq\sqrt{P_{t}},\leavevmode\nobreak\ \forall t\in\mathcal{T}$ (13c)
where $\widetilde{\mathbf{F}}_{t}=\text{diag}\left(\left\lVert\mathbf{f}_{1,t}\right\rVert,\dots,\left\lVert\mathbf{f}_{L,t}\right\rVert\right)$. This problem can be solved using the convex-concave procedure outlined in [3].
@@ III-B RIS phase-shift optimization
We now move on to the second part of the AO, where we optimize $\boldsymbol{\theta}$ when the power variables are fixed. The SCNR can be reformulated as $\overline{\textrm{SCNR}}=\boldsymbol{\theta}^{\mathsf{H}}\mathbf{Q}\boldsymbol{\theta}$, where
@@
We have used $\text{Tr}(\mathbf{A}\mathbf{X}\mathbf{B}\mathbf{X}^{\mathsf{H}})=\text{vec}^{\mathsf{H}}(\mathbf{X})\left(\mathbf{B}^{\top}\otimes\mathbf{A}\right)\text{vec}(\mathbf{X})$ and $\text{vec}^{\mathsf{H}}(\text{diag}(\mathbf{x}))(\mathbf{A}\otimes\mathbf{B})\text{vec}(\text{diag}(\mathbf{x}))=\mathbf{x}^{\mathsf{H}}(\mathbf{A}\odot\mathbf{B})\mathbf{x}$ from [10]. The RIS phase-shift optimization problem under consideration is still non-convex: We then employ the minorization-maximization (MM) algorithm to maximize a convex lower bound on the objective function [11] around the local point $\boldsymbol{\theta}^{(s)}$. We can now define the second subproblem of the AO optimization as
@@
where $\alpha$ is an auxiliary variable and $\mathbf{Q}^{(+)}$ and $\mathbf{Q}^{(-)}$ represent the positive and negative semidefinite parts of $\mathbf{Q}$. Note that we have relaxed the unit modulus constraint, as it is observed that the solution satisfies this constraint with equality anyway. This is a convex problem and can be solved at each iteration until convergence is achieved. The steps of the AO algorithm are outlined in Algorithm 1.
@@ IV GLRT Detector
One of the main metrics used to assess the sensing performance is the probability of correct detection, implemented through a generalized likelihood ratio test (GLRT) by extending the detector in [7] to a cell-free massive MIMO system based on the optimization procedure developed in the previous section. We operate under the assumption that the transmit signals are known by the receiving APs, thanks to the cell-free architecture. We define the vector containing the combined sensing observation of all receiving APs during timeslot $\iota$ as
$\mathbf{y}[\iota]=[\mathbf{y}_{1}^{\top}[\iota],\dots,\mathbf{y}_{R}^{\top}[\iota]]^{\top}=\mathbf{V}[\iota]\boldsymbol{\xi}+\mathbf{U}\mathbf{w}[\iota]+\mathbf{n}[\iota],$
$\mathbf{V}[\iota]=\textrm{blkdiag}(\mathbf{V}_{1}[\iota],\dots,\mathbf{V}_{R}[\iota])\in\mathbb{C}^{MR\times TR},$
$\mathbf{V}_{r}[\iota]=\left[\mathbf{E}_{1,r}\mathbf{s}_{1}[\iota],\dots,\mathbf{E}_{T,r}\mathbf{s}_{T}[\iota]\right]\in\mathbb{C}^{M\times T},$
$\mathbf{n}[\iota]=[\mathbf{n}_{1}^{\top}[\iota],\dots,\mathbf{n}_{R}^{\top}[\iota]]^{\top},\leavevmode\nobreak\ \boldsymbol{\xi}=[\xi_{1,1},\dots,\xi_{T,R}]^{\top},$
$\mathbf{U}=\textrm{blkdiag}\left(\mathbf{U}_{1},\dots,\mathbf{U}_{R}\right),\mathbf{w}[\iota]=[\mathbf{w}_{1}^{\top}[\iota],\dots,\mathbf{w}_{R}^{\top}[\iota]]^{\top}.$
@@
where $\overline{\mathbf{V}}$ is computed as $\mathbf{V}$ by using $\overline{\mathbf{E}}_{t,r}$. The adoption of a GLRT method is motivated by the unknown distribution of $\mathbf{w}[\iota]$, however, we know that $\mathbf{y}_{\mathcal{H}_{0}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{MR})$ and $\mathbf{y}_{\mathcal{H}_{1}}[\iota]-\mathbf{U}\mathbf{w}[\iota]\sim\mathcal{CN}(\mathbf{0},\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})$, where $\mathbf{R}[\iota]=\mathbb{E}\left\{\overline{\mathbf{V}}[\iota]\boldsymbol{\xi}\boldsymbol{\xi}^{\mathsf{H}}\overline{\mathbf{V}}^{\mathsf{H}}[\iota]\right\}=\overline{\mathbf{V}}[\iota]\text{diag}(\delta^{2}_{1,1},\dots,\delta^{2}_{T,R})\overline{\mathbf{V}}^{\mathsf{H}}[\iota]$. Following the derivation in [7], the GLRT test can be rewritten as
$\sum_{\iota=1}^{\tau}\mathbf{y}^{\mathsf{H}}[\iota]\Big{(}\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{(}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]\mathbf{U}\Big{)}^{-1}\mathbf{U}^{\mathsf{H}}\boldsymbol{\Xi}[\iota]+$
$\sigma^{-2}(\mathbf{I}_{MR}-\mathbf{U}\mathbf{U}^{\mathsf{H}})-\boldsymbol{\Xi}[\iota]\Big{)}\mathbf{y}[\iota]\lessgtr^{\mathcal{H}_{0}}_{\mathcal{H}_{1}}\widetilde{\lambda}_{d}$ (17)
where $\boldsymbol{\Xi}[\iota]=(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR})^{-1}$. $\widetilde{\lambda}_{d}=\ln(\lambda_{d})-\ln\left(\det\left(\sigma^{2}\mathbf{I}_{MR}\right)^{\tau}/{\prod_{\iota=1}^{\tau}\det\left(\mathbf{R}[\iota]+\sigma^{2}\mathbf{I}_{MR}\right)}\right)$ is a revised threshold whose value should be chosen according to the desired false alarm probability.
@@ V Numerical Results
Figure 3: Probability of detection vs RCS power with CNR = $20$ dB.
[…]
Figure 4: Probability of detectionvs the RCS power with a CNR=$40$ dB.
@@ V-A Clutter awareness and cluster number’s impact
Fig. 3 shows this probability of detection vs. the RCS power, where we assumed that $\delta_{1,1}^{2}=\dots=\delta_{T,R}^{2}=\delta^{2}$ for a CNR of $20$ dB. We can see that a clutter-aware system outperforms a clutter-unaware one: The clutter-unaware detector is defined as $\mathbf{T}[\iota]=\sigma^{-2}\mathbf{I}_{MR}-\boldsymbol{\Xi}[\iota]$. When $C_{2}=C_{3}=C_{4}=20$, denoted by $C_{234}$, the LoS assumption is no longer realistic, thus generating a performance decrease. Our optimized model greatly outperforms the benchmark approach, consisting of a random RIS phase-shift matrix and an equal power allocation between the beams. Fig. 4 shows what happens when the CNR rises to $40$ dB. We see that clutter awareness rises in importance, widening the gap between the clutter-aware and unaware detectors.
@@
Figure 5: Impact of a dedicated sensing stream onto the probability of detection with a CNR=$20$ dB.
@@ V-B Do we need dedicated sensing symbols?
We have analyzed a system without sensing-dedicated precoding. Fig. 5 compares the detection performance of systems with and without dedicated sensing precoding, both in the presence and absence of clutter awareness. We can immediately see that removing the sensing streams implies a performance degradation. Interestingly, the system without sensing dedicated precoding seems to suffer less from the lack of clutter awareness at a low RCS value.
@@ VI Conclusions
We have investigated a cell-free massive MIMO ISAC network jointly serving multiple communication UEs while detecting a potential target through an RIS in a multi-static fashion. To this end, the transmit power allocation and the RIS phase-shift matrix are optimized to increase the target detection probability. Numerical results show that clutter awareness plays a crucial role in the detection performance and that including a dedicated sensing stream benefits the system. The mmWave channel structure allows the system to assume LoS channels and still achieve good performances when the channel is composed of a high number of clusters concentrated in a small angular range.